Forty Is 40% Less Than What Number?
Answer by benazir.sj@gmail.com(109)

Forty Is 40% Less Than What Number?
solution :
40% less means 100-40 =60%
let the number be x
now the eqaution becomes
60%of x=40
60/100*x=40
6x/10=40
6x=40*10
x=400/6
x=66.66
The number is 66.66